OpenStudy (anonymous):

What is the value of the expression w – 4 when w = –17?

OpenStudy (anonymous):

Can you type out the expression? It didn't come through D:

OpenStudy (anonymous):

Nevermind i derped. It's just -17-4 as the equation. Can you type this into your calculator?

OpenStudy (anonymous):

the awser wil be w = 5/4 and w = -4

OpenStudy (anonymous):

I did not understand what you said @iamcaleb19 and @ohlookidontunderstandanotherthing what do I type in my calculator

OpenStudy (anonymous):

You type -17-4. This is just a substitution problem.

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

Np! :D

OpenStudy (anonymous):

this is whar it look like by the way and thanks @ohlookidontunderstandanotherthing

OpenStudy (anonymous):

heres the answer: w = 5/4 and w = -4

OpenStudy (anonymous):

ty

OpenStudy (anonymous):

got to go

OpenStudy (anonymous):

Is the answer -4?